OpenStudy (anonymous):

Simplify the following. (2x/(x^2-6x+9)) - (1/(x+1)) - (8/(x^2-2x-3))

OpenStudy (anonymous):

\[\frac{ 2x }{ x^2 - 6x +9 } - \frac{ 1 }{ x + 1 } - \frac{ 8 }{ x^2 - 2x - 3 }\]

OpenStudy (anonymous):

So let's do one step at a time. I would recommend you factor them all out first.